OpenStudy (anonymous):

Find the value of Q in the following system so that the solution to the system is {(x,y) : x-3y=4} x-3y=4, 2x+qy=8

OpenStudy (anonymous):

Q = -6

OpenStudy (anonymous):

1) First eqation x-3y=4 2) 2x+qy=8 It looks like if first equations is multiplied by 2 we get second equation, So let do it 2x-6y=8 So we see that in place of q is -6

OpenStudy (anonymous):

Where do you see a,b,or c

OpenStudy (anonymous):

x = 4+ 3y Thus... 2(4 +3y) + qy = 8 8 + 6y + qy = 8 6y +qy = 0 qy = -6y q = -6

OpenStudy (anonymous):

i dont that one girl said it the way it is supposed to be written but everytime i do it it gets messed up is like this Find the value of Q in the following system so that the solution to the system is {(x,y) : x-3y=4} x-3y=4 2x+qy=8

OpenStudy (anonymous):

{(x,y) : x-3y=4} x-3y=4 2x+qy=8

OpenStudy (anonymous):

Kassia did that help?

OpenStudy (anonymous):

yeah so the answer is definetly -6?